OpenStudy (anonymous):

ball1, with a mass of 20kg, is moving to the right at 20 m/s. At what veloctiy should ball 2, with a mass of 40 kg, move so that they both come to a standstill upon collision?

OpenStudy (anonymous):

Actually more suited for Physics Conservation of momentum m1 v1 + m2 v2 = 0 (20kg)(20 m/s) +(40 kg)(v2) = 0 v2 = -400/40 = -10 m/s